Suppose Lester’s Lassos produced 250 lassos and sold each for $100. Lester’s total revenue was _____ dollars.

Suppose Sam’s Sandal Co. makes only one kind of sandal, which it sells for $50 a pair. If Sam’s sells 500 pairs of sandals, then its total revenue will be _____.

Suppose Sam’s Sandal Co. makes only one kind of sandal, which it sells for $50 a pair. If Sam’s sold 500 pairs of shoes and had total cost of $10,000, then Sam’s profit would be _____.
Answer #1

Answer

Suppose Lester’s Lassos produced 250 lassos and sold each for $100. Lester’s total revenue was _____ dollars.

Total revenue =price * quantity

=250*100

=$25000

Suppose Sam’s Sandal Co. makes only one kind of sandal, which it sells for $50 a pair. If Sam’s sells 500 pairs of sandals, then its total revenue will be _____.

Total revenue =price * quantity

=50*500

=$25000

Suppose Sam’s Sandal Co. makes only one kind of sandal, which it sells for $50 a pair. If Sam’s sold 500 pairs of shoes and had total cost of $10,000, then Sam’s profit would be _____.

Profit =total revenue -total cost

TR=50*500

=25000

Profit=25000-10000

=$15000

The profit is $15000